OpenStudy (anonymous):

R = 60, L = 2, C = .0025, E(t) = 100e^-t, I(0) = 0, Q(0) = 0. Find I(t).

OpenStudy (anonymous):

LI'(0) + RI(0) + (1/C)*Q(0) = E(0) 2I'(0) + 60I(0) + 400*Q(0) = 100

OpenStudy (anonymous):

2I'(0) = 100 -> I'(0) = 50

OpenStudy (anonymous):

LI'' + RI' + (1/C)*I = E'(t) 2I'' + 60I' + 400I = E'(t)

OpenStudy (anonymous):

2I'' + 60I' + 400I = -100e^-t 2m^2+60m+400=0 m^2+30m+200=0 (m+10)(m+20)=0 m=-10 , -20 Yc=C1e^-10x + C2e^-20x try to solve for Yp=Ae^x

OpenStudy (anonymous):

if Yp=Ae^x what is Yp'=____? and Yp''=____?

OpenStudy (anonymous):

@ amaes

OpenStudy (anonymous):

so that Y(t)= Yc+Yp